OpenStudy (anonymous):

Find three consecutive even integers, such that 7 times the middle integer is 10 more than the sum of first and third. Please an explanation would be nice.

OpenStudy (anonymous):

Nope sorry

OpenStudy (anonymous):

7y=10+(x+z) \[7*4=10+(2+6)=18\] \[28\neq 18\]

OpenStudy (anonymous):

so let your integers be x,(x+1),(x+2).. then form an equation.. here the middle integer is (x+1) so 7 times of it wud be...&(x+1)... it is now given that 7 times the mid integer is 10 more than sum of first and third.. so equation is 7(x+1)-(x+(x+2)=10

OpenStudy (anonymous):

Can you solve it now?? @ccardenas2322 ??

OpenStudy (anonymous):

Okay let me try :)

OpenStudy (anonymous):

keep me posted..

OpenStudy (anonymous):

if x =2 , the middle integer is then odd

OpenStudy (anonymous):

i believe it's on the right track but you need to make it x (x+2) (x+4)

OpenStudy (anonymous):

yeah sorry it shud be x,(x+2),(x+4)

OpenStudy (anonymous):

so its 7(x+2)-x(x+4)=10

OpenStudy (anonymous):

no its 7(x+2+-{x+(x+4)}=10

OpenStudy (anonymous):

That equation looks confusing,but i'll try.

OpenStudy (anonymous):

\[7(x+2)=10+((x+(x+4))\] \[7x+14=10+(2x+4)\] \[7x+14=10+2x+4\] \[5x=-4+4\] \[x=0\]

OpenStudy (anonymous):

does this make sense?

OpenStudy (anonymous):

however im not sure if 0 is even? lol

OpenStudy (anonymous):

okay that one makes more sense so the integers are 0,2,4

OpenStudy (anonymous):

yeah i guess you're correct.

OpenStudy (anonymous):

well check to see =] using the equality if you let x =0 ,y=2, z=4 \[7y=10+(x+z)\]

OpenStudy (anonymous):

well it does satisfy your condition..

OpenStudy (anonymous):

It does thank you guys ! :)

OpenStudy (anonymous):

sure thing..

OpenStudy (anonymous):

thanks for the medal @Outkast3r09 ..